OpenStudy (anonymous):

The position of a particle moving in the xy plane is given by r⃗ =2.0cos(3.0rad/st)i^+2.0sin(3.0rad/st)j^, where r is in meters and t is in seconds. Show that this represents circular motion of radius 2.0 m centered at the origin.

OpenStudy (agent0smith):

Write it like parametric equations, like this (also, don't write rad/s etc in there, makes it look way more confusing) x = 2cos3t y = 2sin3t from those you get x/2 = cos3t and y/2 = sin3t Now use the fact that \[\Large \cos^2 \theta +\sin^2 \theta = 1\]

OpenStudy (agent0smith):

You will want to square both sides of each of these: x/2 = cos3t and y/2 = sin3t

OpenStudy (anonymous):

Thank you. That is actually how we were given the question in MasteringPhysics. That have been where some of my confusion came from.